OpenStudy (anonymous):

Design five equations, like 3x + 4 = 16. Make sure that the solutions to the equations are integers. In other words, 3x + 4 = 15 wouldn't be acceptable, because the solution would be x = 11/3.

OpenStudy (anonymous):

3x + 4 = 13 3x + 4 = -5 3x + 4 = 22 3x + 4 = -8 3x + 4 = 31

OpenStudy (anonymous):

Those solutions don't work.

OpenStudy (anonymous):

They must have integer solutions.

OpenStudy (anonymous):

Integers.

OpenStudy (anonymous):

Ah wait, yes they do, I misread the question.

OpenStudy (anonymous):

ummm they do work

OpenStudy (anonymous):

Yes, I misread the question. Good job. :D

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

thank u sir

OpenStudy (anonymous):

Yes it does.

OpenStudy (anonymous):

3x + 4 = 31 3x = 27 x = 9

OpenStudy (anonymous):

Give him a medal.

OpenStudy (anonymous):

Miss read it also ha ha it does